RC meeting and Run Plan, Monday Oct. 2, 2023
Return to the RC Meeting Calendar
- 1 RC Meeting Information
- 2 Useful Links
- 3 Procedures
- 4 Status Update
RC Meeting Information
Time: Monday - Friday at 3:30pm
- Location: Counting house conference room
- ZoomGov Meeting Connection Information
Hall A call-in instructions:
- Call RC first.
- Techs on site 7am-4:30pm Mon-Thu, 7-11am Fri. RC will call Hall A work coordinator (Zach) during normal working hours. After hours, RC calls Tech on-call.
- After 9pm everything non-emergency needs to be deferred. What is emergency and what's not? Call RC. RC talks with Mark to decide.
Hall Access procedure:
- Access check list (binder on the counting room desk) needs to be filled out before making an access.
- Scan (take pictures with phone) and post it to the logbook.
- Any work inside the SBS, BB boundary need to have PS locked out -- requires two QEWs to throw the big switch. Exception: field measurements.
New instruction for shift crew when changing target:
- Beam restoration instructions have been updated. Printed version is in the counting room.
- When we change target, make sure the SOFs of Target BPMs are set to the correct values.
- SOFs for different targets: Needs an update after restoring beam tonight (Note: SOFs for major target positions are on the Control Room whiteboard.)
Beam was finally restored at the end of Saturday swing shift, and we have been running production on the PolHe3 cell until 7:30 this morning. Yes, 8 hours was taken with the wrong BBCAL HV values, but that was rectified and we are now using the correct settings. At half past 7 this morning, the oven temperature started to drop, so the lasers were turned off. After the MCC meeting this morning, Bill undertook an escorted access to investigate the issue. From this access, it was found that the problem was going to be more involved than originally thought; a controlled access was organized so that this could be fixed. That access is now complete: the issue is likely to have been from a fuse in the box that the variacs run through; a fuse which has been replaced. The oven is back on, and we are now only waiting for the beam to be restored.
- Beam has been down since just after 9am this morning, for MCC tests to get things ready for Hall B to start running during swing shift.
- 4-pass separator
- No maintenance day on the MCC white board (until 10/10/2023).
Data collected :
- Accumulated corrected charge collected: 1.884C since last Wednesday (plus 0.569C at the wrong BBCAL HV settings).
- 1.129C of 'good' charge on Sunday (10/01)
- 0.756C today (10/02 - all in OWL shift)
Hall A Subsystems Status
- The oven temperature started dropping seriously at 7:30 this morning. The lasers were turned off, so that the hall could be accessed to check it out.
- It looks like the issue was with a fuse that the variacs were going through (which needed a controlled access anyway, so no time was wasted). This has been fixed, and the oven is now warming up.
- I went into the hall to check the hodoscope HV to try to fix the issue with 48R and 49R. The HV crate has been powercycled, and we shall see if this works when we are back in production.
Hall access requests
- If anyone wants to make an access, let me know! There is no planned maintenance day before Tuesday 10th, so any accesses need to be opportunistic.
- Still need to fill open slots
- 10/5 - need a TO for OWL
- Please check when your shifts are!
Please sign up for any open shifts!
Short Term Run plan
- Beam recovery to the hall
- Production run on polarized He3 target @45uA(!) beam current.
- If the target polarization drops dramatically during 45uA running, lower the beam current to 40uA.
==> List of items in case He3 production running isn’t ready:
- Take data on reference cell at 10 atm H2 (portions that add up to 1 day) [85% (4200 elastic events) more data required for Bigbite momentum calibration]
- BigBite Sieve slit => only during day (Techs needed)
- with 8 carbon foils and take optics for 1 hr at 40 uA (Sean/Andrew - confirm time)
- 1 hour for single foil (magnets off) => Chance BBcal HV setting to appropriate values
- BigBite/single foil location reconstruction (~20 mins?) - BigBite + HCAL detector experts on call consulting analysis experts like Andrew/Sean (magnet on)